# Interpretable Multi-task Learning with Shared Variable Embeddings

### v1

@@ Abstract
This paper proposes a general interpretable predictive system with shared information. The system is able to perform predictions in a multi-task setting where distinct tasks are not bound to have the same input/output structure. Embeddings of input and output variables in a common space are obtained, where the input embeddings are produced through attending to a set of shared embeddings, reused across tasks. All the embeddings are treated as model parameters and learned. Specific restrictions on the space of shared embedings and the sparsity of the attention mechanism are considered. Experiments show that the introduction of shared embeddings does not deteriorate the results obtained from a vanilla variable embeddings method. We run a number of further ablations. Inducing sparsity in the attention mechanism leads to both an increase in accuracy and a significant decrease in the number of training steps required. Shared embeddings provide a measure of interpretability in terms of both a qualitative assessment and the ability to map specific shared embeddings to pre-defined concepts that are not tailored to the considered model. There seems to be a trade-off between accuracy and interpretability. The basic shared embeddings method favors interpretability, whereas the sparse attention method promotes accuracy. The results lead to the conclusion that variable embedding methods may be extended with shared information to provide increased interpretability and accuracy.
@@ 1 Introduction
The ability to extract common information from varied settings has long been one of the central challenges in machine learning. The degree to which the considered domains differ and the complexity of the domains themselves has grown considerably over time. Artificial neurons [24] linked in a physical perceptron model [34] were used to distinguish, through a weight update procedure, the side on which a punch card had been marked. CNNs [13, 21] are able to identify similar patterns in distinct areas of an image. Word embeddings [4, 28] have been used to obtain representations fusing information from different contexts. Reinforcement learning agents achieved relatively high performance in a number of Atari games without changes to the model architecture [30]. Generative methods, such as diffusion models [33] produce high-fidelity imagery based on information from a provided text prompt.
Multi-task learning (MTL) is a specific area of machine learning concerned with approaches that attempt to simultaneously solve more than one task [6, 7, 8, 35, 9]. With the evolution of deep learning architectures over the years, we have seen a sharp increase in the interest in MTL, e.g. in hard parameter sharing methods [19, 11], soft parameter sharing methods [29, 14], decoder models [5, 39]. MTL also has strong links to self-supervised learning (SSL) — a learning paradigm where models can be pretrained on unlabelled data in order to improve their performance or data efficiency on downstream tasks [28, 10, 40, 3, 1].
MTL and SSL can be interpreted as settings in which knowledge about one task facilitates the learning of another one. This is usually done via a choice of tasks that share significant structure, e.g. semantic segmentation, human parsing, monocular depth estimation, etc. This is in stark contrast to real-world prediction scenarios, where typically no a priori structure is given. The difficulty in obtaining meaningful structure from unrelated tasks has led research on MTL to mostly focus on related tasks, even though there is a body of work suggesting that solving not obviously related tasks may actually be helpful in making sound predictions [22, 26].
@@ 2 Background
The notion of word embeddings [4] can be extended to variable embeddings (VEs) [27] by treating the $i$-th variable as being associated with two elements:
@@
As far as specific choices of architectures of the encoder and decoders are concerned, we follow the setup presented in [27] where the conditioning on variable embeddings is done via FiLM layers [31].
The described procedure shows specific advantages, e.g. the possibility to handle tasks with different dimensions of input and output spaces, the ability to recover structure on small-scale problems, and relatively good performance on a range of tasks. On the flip side, it does not reuse the obtained embeddings between variables and it does not lend itself readily to interpretation for real-world classification datasets. In principle, a variable embedding is obtained for each observed and target variable, so the embeddings can be compared in their common space or projected into a lower-dimension space for visualization using methods such as t-SNE [18] or UMAP [25]. In reality, however, this turns out to be problematic for more complex data. For instance, for the real world dataset of UCI-121 [12, 20], the vanilla variable embeddings approach produces embeddings which seem to differentiate between the observed variables, common target variables and uncommon target variables [27], but we do not have any more information on the relations between the variables themselves.
@@ 3 Method
Figure 1: The overview of the shared variable embeddings method. The variable space contains both the observed and target variables which are associated with their learnable variable embeddings (VEs). The observable variables are first linked to raw VEs which are used as queries in the attention mechanism. A separate set of shared VEs plays the role of both keys and values. The processed VEs are the output of attention. Together with the corresponding variable values they are processed, each (value, VE) pair separately, by the encoder. The outputs of the encoder are summed and passed to the initial decoder. The target variable of interest is directly linked to its VE and this VE is passed with the output of the initial decoder to the final decoder to actually perform the prediction of the value of the target variable of interest. Additional details of the architecture are available in the Appendix in Figure 4.
@@ 3.1 Shared variable embeddings
Let us consider $N$ observed variables and a set of $D$ shared embeddings $\{\mathbf{s}_{k}\}_{k=1}^{D}$, with $\mathbf{s}_{k}\in\mathbb{R}^{C}$. The associated shared embedding matrix is $\mathbf{S}_{D\times C}$, where $C$ is the dimension of both the embedding space of observed variables and the shared embedding space. In order to enforce the reuse of information between variable embeddings, we would like $D\ll N$. We relate the raw (initial) variable embeddings to the shared embeddings via attention.
For the raw variable embedding matrix $\mathbf{Z}_{N\times C}$, we follow the standard attention procedure [38]:
@@ 3.2 Training
where $t_{j}$ is a +1/1+1/-1+ 1 / - 1 encoding of the actual target and $\hat{y}_{j}$ is the prediction of the value of the $j$-th target variable for the given task obtained from the encoder/decoder architecture with shared variable embeddings as in Eq. 6.
@@ 3.3 Imposing independence of shared variable embeddings through additional structure
One question that can be asked of the shared embedding matrix $\mathbf{S}_{D\times C}$ is that of structure. In particular, it could be argued that the learning process does not explicitly require the shared embeddings $\{\mathbf{s}_{k}\}_{k=1}^{D}$ to be independent from one another. We consider different notions of independence and several approaches to encourage it in the shared embeddings.
3.3.1 Orthogonalization
For a simple notion of independence, we consider an embedding independent from other embeddings when it is not a linear combination of them. With this, $r_{\mathbf{S}}=\text{rank}(\mathbf{S})$ is a measure of independence. In a realistic setting, we might still have $r_{\mathbf{S}}=C$ even if multiple embeddings are approximately linearly dependent. An operational measure of the rank of $\mathbf{S}$ would require to address this drawback and we describe such a measure $q_{\mathbf{S}}$ in Section 3.3.2. The results from Section 4.2 show that the proposed training procedure results in $q_{\mathbf{S}}<C$. A straightforward way to build in more independence is to require $\mathbf{S}$ to consist of orthonormal vectors. With an additional assumption of $D=C$, this would translate into an orthogonality requirement, which could be incorporated in the loss function:
3.3.2 Stable rank
Instead of focusing on restricting $\mathbf{S}$, it is possible to explicitly add $r_{\mathbf{S}}$ to the loss function. A significant drawback of this is the discontinuous characteristic of the rank measure, which makes it unsuitable for gradient descent. To address this, we rely on a continuous proxy. Let us consider a matrix $\mathbf{A}_{N\times M}$ with $\sigma_{i}(\mathbf{A})$ being its $i$-th singular value. The Frobenius norm of $\mathbf{A}$ is defined as $\left\lVert\mathbf{A}\right\rVert_{F}^{2}=\text{tr}(\mathbf{A}\mathbf{A}^{T})=\sum_{i,j}{\mathbf{A}_{i,j}^{2}}=\sum_{i}{\sigma_{i}^{2}}$. The stable rank [37] of $\mathbf{A}$ is then defined as:
and $\text{sr}(\mathbf{A})\leq\text{rank}(\mathbf{A})$.
3.3.3 Von Neumann entropy
$V(\mathbf{A})=-\sum_{i}{\sigma_{i}^{2}\ln{\sigma_{i}^{2}}}$ (11)
3.4 Sparse attention
In a procedure orthogonal to inducing structure in the shared embedding matrix, one can also restrict the way in which the actual shared embeddings are combined to form the processed embeddings. One drawback of the standard attention mechanism is that it assigns non-zero weights to all the value vectors. This means that even components with marginal similarity to the keys are present in the final linear combinations. A potential solution would be to make the output of the attention mechanism not rely on values with small similarity scores. In order to keep the whole mechanism differentiable, we adopt the $\alpha$-entmax method [32]. Let us denote the $d$-probability simplex by $\triangle^{d}=\{\mathbf{p}\in\mathbb{R}^{d}:\mathbf{p}\geq 0,\left\lVert p\right\rVert_{1}=1\}$. Sparsemax [23] is defined as:
A family of Tsallis $\alpha$-entropies [36] can be defined for $\alpha\geq 1$ as:
where $H^{S}(\mathbf{p})=-\sum_{j}{p_{j}\ln{p_{j}}}$.
4 Experiments
We validate the ability of shared variable embeddings to solve real-life classification tasks and to help in interpretability on the UCI-121 [12, 20] dataset. In the experiments, we use the hyperparameter values and the learning setup from [27]. For the shared variable embeddings, we choose $D=C=128$. We provide quantitative comparisons of the proposed method against a strong baseline: the variable embedding method without shared embeddings. We also consider the results with restrictions on the shared embedding matrix and on the attention mechanism, as proposed in Section 3. Additionally, we provide qualitative comparisons for the basic version of our method and for its configurations with constraints on the shared embedding matrix and with sparse attention.
4.1 Classification capability
Results in terms of best test set accuracy are presented in Table 1. We use the vanilla variable embedding method without shared embeddings [27] as a strong baseline. Quantitative assessment shows that the shared embedding approach is able to achieve a classification accuracy in the range of the results from the baseline. At the same time, the $1.05$-entmax sparse attention method is able to surpass the accuracy levels of both the baseline and the shared embedding method with full attention. Notably, the baseline requires additional fine-tuning on each of the 121 datasets while our approaches do not.
Details of the training process are displayed in Figure 2. Both the shared embedding method and the $1.05$-entmax method show similar characteristics through the earlier part of the training process. A major difference, however, is that the entmax method hits the stop criterion significantly earlier and provides a higher final test set accuracy. The stable rank method with $\alpha_{\text{sr}}=0.05$ hits visibly lower accuracy levels throughout training, while requiring more steps than the $1.05$-entmax method. Figure 3(a) directly compares the number of steps before reaching the maximum test set accuracy. In particular, the $1.05$-entmax model reaches its peak test set accuracy after $10,300$ steps compared to $16,200$ for the shared embedding method, which is a $36.4\%$ decrease in training time measured in steps. The stable rank model with $\alpha_{\text{sr}}=0.05$ requires $10,600$ steps.
Ablation studies for methods involving orthogonalization, stable rank and von Neumann entropy as means to enforce independence in the shared embeddings are presented in Table 2. These results suggest that orthogonalization and, in particular, von Neumann entropy have an adverse effect on the final classification accuracy, while the stable rank restrictions do not seem to improve the results relative to straightforward shared embeddings but they do not decisively hurt them either.
001111222233334444555566667777888899991010101011111111121212121313131314141414151515151616161617171717181818180.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91.01.01.01.0training step (k)test accuracySESRENT
Figure 2: UCI-121 test set accuracy for a given train step (in thousands). SE - shared embedding method, ENT - $1.05$-entmax with embeddings initialized from $\mathcal{N}(0,1)$, SR - stable rank with $\alpha_{\text{sr}}=0.05$.
An extensive ablation for the sparse attention methods is presented in Table 3. The $\alpha$-entmax approaches are evaluated in two distinct settings. In the first one, $\alpha$ is picked as a hyperparameter, constant across the whole training procedure. In the second one, $\alpha$ is treated as a model parameter, with an initial value, and is optimized with gradient descent. In this case, the final optimized value of $\alpha$ is reported. The evaluations of the $\alpha$-entmax methods show that it is important to adjust the weight initialization procedure of the model. For embeddings initialized as in the vanilla variable embedding approach, the final accuracy suffers. Increasing the standard deviation of the normal distribution from which the initial weights are sampled markedly improves the test accuracy. In particular, for the $1.05$-entmax method with a standard normal distribution used for initialization, the accuracy reaches levels higher than for any other considered setup, with a significantly shortened training time. This suggests that for the particular tackled set of tasks, an attention mechanism with moderate induced sparsity provides slight advantages in terms of accuracy and significant advantages in terms of training time relative to the standard attention mechanism.
SESRENT0055551010101015151515
RANDSEENTSR0020202020404040406060606080808080100100100100120120120120
Figure 3: (a) Training steps needed to reach best test set accuracy. (b) Stable rank of the shared embedding matrix after training - best accuracy model. SE - shared embedding method, ENT - $1.05$-entmax with embeddings initialized from $\mathcal{N}(0,1)$, SR - stable rank with $\alpha_{\text{sr}}=0.05$, RAND - random embedding matrix with entries from $\mathcal{N}(0,1)$.

### v2

# My Publication Title — Multiple Authors

@@ Abstract
This paper proposes a general interpretable predictive system with shared information. The system is able to perform predictions in a multi-task setting where distinct tasks are not bound to have the same input/output structure. Embeddings of input and output variables in a common space are obtained, where the input embeddings are produced through attending to a set of shared embeddings, reused across tasks. All the embeddings are treated as model parameters and learned. The approach is distinct from existing vector quantization methods. Specific restrictions on the space of shared embedings and the sparsity of the attention mechanism are considered. Experiments show that the introduction of shared embeddings does not deteriorate the results obtained from a vanilla variable embeddings method. We run a number of further ablations. Inducing sparsity in the attention mechanism leads to both an increase in accuracy and a significant decrease in the number of training steps required. Shared embeddings provide a measure of interpretability in terms of both a qualitative assessment and the ability to map specific shared embeddings to pre-defined concepts that are not tailored to the considered model. There seems to be a trade-off between accuracy and interpretability. The basic shared embeddings method favors interpretability, whereas the sparse attention method promotes accuracy. The results lead to the conclusion that variable embedding methods may be extended with shared information to provide increased interpretability and accuracy.
@@ 1 Introduction
The ability to extract common information from varied settings has long been one of the central challenges in machine learning. The degree to which the considered domains differ and the complexity of the domains themselves has grown considerably over time. Artificial neurons (Mcculloch and Pitts 1943) linked in a physical perceptron model (Rosenblatt 1958) were used to distinguish, through a weight update procedure, the side on which a punch card had been marked. CNNs (Fukushima 1980; LeCun et al. 1989) are able to identify similar patterns in distinct areas of an image. Word embeddings (Bengio, Ducharme, and Vincent 2000; Mikolov et al. 2013) have been used to obtain representations fusing information from different contexts. Reinforcement learning agents achieved relatively high performance in a number of Atari games without changes to the model architecture (Mnih et al. 2015). Generative methods, such as diffusion models (Rombach et al. 2022) produce high-fidelity imagery based on information from a provided text prompt.
Multi-task learning (MTL) is a specific area of machine learning concerned with approaches that attempt to simultaneously solve more than one task (Caruana 1993, 1994, 1996; Thrun and O’Sullivan 1996; Caruana 1997). With the evolution of deep learning architectures over the years, we have seen a sharp increase in the interest in MTL, e.g. in hard parameter sharing methods (Hu and Singh 2021; Cui et al. 2021), soft parameter sharing methods (Misra et al. 2016; Gao et al. 2019), decoder models (Brüggemann et al. 2021; Ye and Xu 2023). MTL also has strong links to self-supervised learning (SSL) — a learning paradigm where models can be pretrained on unlabelled data in order to improve their performance or data efficiency on downstream tasks (Mikolov et al. 2013; Chen et al. 2020; Zbontar et al. 2021; Bardes, Ponce, and LeCun 2022; Assran et al. 2023).
MTL and SSL can be interpreted as settings in which knowledge about one task facilitates the learning of another one. This is usually done via a choice of tasks that share significant structure, e.g. semantic segmentation, human parsing, monocular depth estimation, etc. This is in stark contrast to real-world prediction scenarios, where typically no a priori structure is given. The difficulty in obtaining meaningful structure from unrelated tasks has led research on MTL to mostly focus on related tasks, even though there is a body of work suggesting that solving not obviously related tasks may actually be helpful in making sound predictions (Mahmud and Ray 2007; Meyerson and Miikkulainen 2019).
@@ 2 Background
The notion of word embeddings (Bengio, Ducharme, and Vincent 2000) can be extended to variable embeddings (VEs) (Meyerson and Miikkulainen 2021) by treating the $i$-th variable as being associated with two elements:
@@
As far as specific choices of architectures of the encoder and decoders are concerned, we follow the setup presented in the Traveling Observer Model (TOM) (Meyerson and Miikkulainen 2021) where the conditioning on variable embeddings is done via FiLM layers (Perez et al. 2018). The general motivation behind VEs is discussed in Appendix L.
The described procedure shows specific advantages, e.g. the possibility to handle tasks with different dimensions of input and output spaces, the ability to recover structure on small-scale problems, and relatively good performance on a range of tasks. On the flip side, it does not reuse the obtained embeddings between variables and it does not lend itself readily to interpretation for real-world classification datasets. In principle, a variable embedding is obtained for each observed and target variable, so the embeddings can be compared in their common space or projected into a lower-dimension space for visualization using methods such as t-SNE (Hinton and Roweis 2002) or UMAP (McInnes, Healy, and Melville 2018). In reality, however, this turns out to be problematic for more complex data. For instance, for the real world dataset of UCI-121 (Fernández-Delgado et al. 2014; Kelly, Longjohn, and Nottingham 2023), the vanilla variable embeddings approach produces embeddings which seem to differentiate between the observed variables, common target variables and uncommon target variables (Meyerson and Miikkulainen 2021), but we do not have any more information on the relations between the variables themselves.
@@ 3 Method
Figure 1: The overview of the shared variable embeddings method. The variable space contains both the observed and target variables which are associated with their learnable variable embeddings (VEs). The observable variables are first linked to raw VEs which are used as queries in the attention mechanism. A separate set of shared VEs plays the role of both keys and values. The processed VEs are the output of attention. Together with the corresponding variable values they are processed, each (value, VE) pair separately, by the encoder. The outputs of the encoder are summed and passed to the initial decoder. The target variable of interest is directly linked to its VE and this VE is passed with the output of the initial decoder to the final decoder to actually perform the prediction of the value of the target variable of interest. Additional details of the architecture are available in Appendix C (Figure 4). The differences between our method and VQ-VAE (van den Oord, Vinyals, and Kavukcuoglu 2017) are highlighted in Appendix I.
@@ 3.1 Shared variable embeddings
Let us consider $N$ observed variables and a set of $D$ shared embeddings $\{\mathbf{s}_{k}\}_{k=1}^{D}$, with $\mathbf{s}_{k}\in\mathbb{R}^{C}$. The associated shared embedding matrix is $\mathbf{S}_{D\times C}$, where $C$ is the dimension of both the embedding space of observed variables and the shared embedding space. In order to enforce the reuse of information between variable embeddings, we would like $D\ll N$. We relate the raw (initial) variable embeddings to the shared embeddings via attention. The increase in model parameters is motivated in Appendix J.
For the raw variable embedding matrix $\mathbf{Z}_{N\times C}$, we follow the standard attention procedure (Vaswani et al. 2017):
@@ 3.2 Training
where $t_{j}$ is a +1/1+1/-1+ 1 / - 1 encoding of the actual target and $\hat{y}_{j}$ is the prediction of the value of the $j$-th target variable for the given task obtained from the encoder/decoder architecture with shared variable embeddings as in Eq. 6. Details of the hinge loss are discussed in Appendix M.
@@ 3.3 Imposing independence of shared variable embeddings through additional structure
One question that can be asked of the shared embedding matrix $\mathbf{S}_{D\times C}$ is that of structure. In particular, it could be argued that the learning process does not explicitly require the shared embeddings $\{\mathbf{s}_{k}\}_{k=1}^{D}$ to be independent from one another. We consider different notions of independence and several approaches to encourage it in the shared embeddings:
  • Orthogonalization: encouraging $\mathbf{S}$ to consist of orthonormal vectors.
  • Stable rank: nudging $\mathbf{S}$ to have high rank.
  • Von Neumann entropy: optimizing for vectors in $\mathbf{S}$ to be independent from the point of view of information theory.
  • Sparse attention: adding sparsity to the attention mechanism.
The specific details of all these approaches are discussed in Appendix A.
@@
4 Experiments
We validate the ability of shared variable embeddings to solve real-life classification tasks and to help in interpretability on the UCI-121 dataset (Fernández-Delgado et al. 2014; Kelly, Longjohn, and Nottingham 2023) (Appendix E). In the experiments, we use the hyperparameter values and the learning setup from (Meyerson and Miikkulainen 2021). Information on code and data availability is included in Appendix N. For the shared variable embeddings, we choose $D=C=128$. We provide quantitative comparisons of the proposed method against a strong baseline: the variable embedding method without shared embeddings. We also consider the results with restrictions on the shared embedding matrix and on the attention mechanism, as proposed in Section 3. We provide qualitative comparisons for the basic version of our method and for its configurations with constraints on the shared embedding matrix and with sparse attention. Additionally, we report the results of extensive ablations (Appendix D), per-task metrics (Appendix F), experiments on additional datasets (Appendix H) and training time (Appendix K).
4.1 Classification capability
Results in terms of best test set accuracy are presented in Table 1. We use the vanilla variable embedding method without shared embeddings (Meyerson and Miikkulainen 2021) as a strong baseline. Quantitative assessment shows that the shared embedding approach is able to achieve a classification accuracy in the range of the results from the baseline. At the same time, the $1.05$-entmax sparse attention method is able to surpass the accuracy levels of both the baseline and the shared embedding method with full attention. Notably, the baseline requires additional fine-tuning on each of the 121 datasets while our approaches do not.
@@ 4.1 Classification capability
Details of the training process are displayed in Figure 2. Both the shared embedding method and the $1.05$-entmax method show similar characteristics through the earlier part of the training process. A major difference, however, is that the entmax method hits the stop criterion significantly earlier and provides a higher final test set accuracy. The stable rank method with $\alpha_{\text{sr}}=0.05$ hits visibly lower accuracy levels throughout training, while requiring more steps than the $1.05$-entmax method. Figure 3(a) directly compares the number of steps before reaching the maximum test set accuracy. In particular, the $1.05$-entmax model reaches its peak test set accuracy after 10 3001030010\,30010 300 steps compared to 16 2001620016\,20016 200 for the shared embedding method, which is a $36.4\%$ decrease in training time measured in steps. The stable rank model with $\alpha_{\text{sr}}=0.05$ requires 10 6001060010\,60010 600 steps.
001111222233334444555566667777888899991010101011111111121212121313131314141414151515151616161617171717181818180.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91.01.01.01.0training step (k)test accuracySVESRENT
Figure 2: UCI-121 test set accuracy for a given train step (in thousands). SVE - shared embedding method, ENT - $1.05$-entmax with embeddings initialized from $\mathcal{N}(0,1)$, SR - stable rank with $\alpha_{\text{sr}}=0.05$.
Ablation studies for methods involving orthogonalization, stable rank and von Neumann entropy as means to enforce independence in the shared embeddings are presented in Table 2. These results suggest that orthogonalization and, in particular, von Neumann entropy have an adverse effect on the final classification accuracy, while the stable rank restrictions do not seem to improve the results relative to straightforward shared embeddings but they do not decisively hurt them either. An extensive ablation for the sparse attention methods is presented in Table 3. The $\alpha$-entmax approaches are evaluated in two distinct settings. In the first one, $\alpha$ is picked as a hyperparameter, constant across the whole training procedure. In the second one, $\alpha$ is treated as a model parameter, with an initial value, and is optimized with gradient descent. In this case, the final optimized value of $\alpha$ is reported. The evaluations of the $\alpha$-entmax methods show that it is important to adjust the weight initialization procedure of the model. For embeddings initialized as in the vanilla variable embedding approach, the final accuracy suffers. Increasing the standard deviation of the normal distribution from which the initial weights are sampled markedly improves the test accuracy. In particular, for the $1.05$-entmax method with a standard normal distribution used for initialization, the accuracy reaches levels higher than for any other considered setup, with a significantly shortened training time. This suggests that for the particular tackled set of tasks, an attention mechanism with moderate induced sparsity provides slight advantages in terms of accuracy and significant advantages in terms of training time relative to the standard attention mechanism.
4.2 Interpretability
SVESRENT0055551010101015151515
RANDSVEENTSR0020202020404040406060606080808080100100100100120120120120
Figure 3: (a) Training steps to reach best test set accuracy. (b) Stable rank of the shared embedding matrix after training - best accuracy model. SVE - shared embedding method, ENT - $1.05$-entmax with embeddings initialized from $\mathcal{N}(0,1)$, SR - stable rank with $\alpha_{\text{sr}}=0.05$, RAND - random embedding matrix with entries from $\mathcal{N}(0,1)$.
  • Compute the attention scores for all raw variable embeddings.
The variables most similar to a shared embedding representing a specific concept would be expected to share some intuitive notion, category or semantic meaning. In our evaluations, we choose $K=5$. To aid in a quantitative as well as qualitative assessment of mapping to concepts, we introduce a measurable and less subjective assignment to real-world concepts in the form of the Subject Area ascribed to a given task/dataset in the UCI repository (Kelly, Longjohn, and Nottingham 2023). Each dataset has one Subject Area (SA) assigned to it from the following $11$ possibilities: Biology (Bio), Business (Bus), Climate and Environment (C&E), Computer Science (CS), Engineering (E), Games (G), Health and Medicine (H&M), Law (L), Physics and Chemistry (P&C), Social Sciences (SS) and Other (O). These categories afford us the option to measure the performance of a given model in terms of interpretability and compare it with other methods.